college students spend $183 more each year on textbooks and course materials than on ccomputer equipment. They spend a total of $819 on textbooks and course materials and computer equpment each year. How much is spent each year on text books and course materials and how much is spent on computer equipment?
Answer by checkley71(8403)

X+(X+183)=819 OR 2X=819-183 OR 2X=636 OR X=636/2 OR X=318 FOR COMPUTERS AND
318+183=501 FOR THE REST.
ADDING THESE TWO WE GET 318+501=819